OpenStudy (anonymous):

Factor. 8x2 + 12x A. 4x(2x + 3) B. 8x(x + 3) C. 8x(x + 4) D. 4x2(2x + 3)

OpenStudy (jdoe0001):

take out common factor \(\bf 8x^2 + 12x\implies ({\color{blue}{ 2\cdot 2}}\cdot 2\cdot {\color{blue}{ x}}\cdot x)+({\color{blue}{ 2\cdot 2}}\cdot 3\cdot {\color{blue}{ x}})\)